Home‎ > ‎Reports‎ > ‎Crew 112 (DOMMEX)‎ > ‎

day01


Crew #112 Log Book for January 28, 2012

Photos of the day





Commander's Report

Date and time: January 27, 2012

Written by Mary Beth Wilhelm Crew 112 Commander

Crew Physical Status: Good.

Time Departed/Returned from EVA: n/a

Summary of Daily Activities:

We arrived at the Hab at 11:00AM and completed a crew changeover with Crew 111b after stopping at the Bull Mountain Market to pick up food study supplies. Crew 111b was very friendly and helpful, and we enjoyed talking to them about their experiences at MDRS and hopes to pursue careers in astronomy in the future. They did an excellent job at cleaning the Hab and the major systems were operational when we arrived. After their departure Crew 112 unpacked, set up the laboratory for our studies, did inventory of food, thoroughly cleaned tables, counters, sink, and bathroom. Additionally, I did a brief orientation tour for the crew members, pointing out the vertebrate fossils and major geologic units. We will have a more detailed lesson tomorrow. I noticed significant scouring tracks from ATV off-roading that were not present when I was here two years ago. Also, we noticed a significant amount of trash (plastic bottles,

Engineering/Hab Maintenance:

Web cams now operational. All other systems operating normally.

Plans for Tomorrow:

AM geology and biology orientation (Hab area)

PM begin transect region scouting for studies (Factory Butte Area)

NorCal Mars Society rover set up and demonstration



Inventory: n/a

Support Requested: some questions sent in email to Jean, some miscellaneous engineering questions (see below)

Please add crew emails to mission support list so we can submit reports

Miscellaneous: BLM manager stopped by today and mentioned that on EVA, we should keep an eye out for a pack of about 100 dogs in the area that have been reported as being aggressive. Future crews should be alerted to this potential hazard.

Health and Safety Report

Date and time: January 28th, 2012, 18:45-07:00

Written by: Noah Warnke, Crew 112 HSO

No incidents. Crew health good all around. I went through the HSO cabinet and pulled out some out-of-date meds, which will be described in the weekly report next week.

Engineering Report

Date and time: 28 January, 2012, 18:00 MDRS
Written by Christopher Haberle, Crew 112: Chief Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): Do I need to be checking these levels? If so where?
Propane Tank (Full/0.75/0.5/0.25/Empty): Do I need to be checking these levels? If so where? Gasoline Tank ATVs (Full/0.75/0.5/0.25/Empty): unchanged

Oil:
ATV Oil Quantity (# of Quarts in storage): N/A
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 22
Trailer Potable Water Tank Level (inches from bottom): 20
Hab Potable Water Tank Level (inches from bottom): 12
Potable Water Meter Reading:

Into GreenHab Meter Reading: 0
Notes/Comments: Where is the Potable Water Meter?

GreenHab: nothing working
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): NO, waiting MDRS lenses
Observatory Notes/Comments: N/A

Transportation:
Hab Car (RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125488.9
Notes/Comments on Hab Car:
Hab car is not at the Hab. We arrived with our NASA government vehicles.

ATV 1 (Opportunity):
Used: No

Oil Checked: No
Fuel Consumed: N/A

Tire Status: Good – N/A
Today's Run Time: 00:00
Season's Run Time: 0:45
Comments: I will begin surveying the vehicles tomorrow.

ATV 2 (Spirit):
Used: NO
Oil Checked: NO
Fuel Consumed: 00%
Tire Status: Good – N/A
Today's Run Time: 00:00
Season's Run Time: 45:00
Comments:

ATV 3 (Viking I):
Used: NO
Oil Checked: NO
Fuel Consumed: 00%
Tire Status: Good
Today's Run Time: 00:00
Season's Run Time: 00:45
Comments:

NorCal Mars Society Rover (Sandstorm):

Used: NO

Number of Batteries Used: 0

Batteries Charged?: 0

Questions/concerns regarding the rover: Tonight I am preparing the rover to be deployed for tomorrow and setting up the base station.

Heating and Ventilation: ok
Thermostat setting upstairs (Farenheit): 72 F
Actual temperature upstairs (Farenheit): 70.2 F

Thermostat setting downstairs (Farenheit): 85 F
Actual temperature downstairs (Farenheit): 52 F
Comments: The Hab was thoroughly cleaned today and the doors were open to air out the smell, hence the discrepancy in temperatures down staris.

Computers/Networking Infrastructure: good.
Notes/comments on internet/computers: Webcams are mostly functional for the time being.

General Engineering Notes/Comments: Nada.